Circular No. 7667

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2001df IN MCG –04-51-5 P. Martin, M. Modjaz, and W. D. Li, University of California at Berke-

ley, report the discovery by LOTOSS (cf. IAUC 7514) of an apparent supernova (mag ∼ 18.0) in unfiltered images taken with the Katzman Automatic Imaging Telescope (KAIT) on June 25.4, July 11.4, and July 18.4 UT. An image taken on June 16.5 under poor conditions also showed a hint of the new object. SN 2001df is located at α = 21^{h}39^{m}18^{s}.15, δ = −25^{o}29′36′′.0 (equinox 2000.0), which is 5′′.0 east and 3′′.4 north of the nucleus of MCG –04-51-5. A KAIT image taken on 2000 Nov. 27.2 showed nothing at this position (limiting mag ∼ 19.0).

T. Matheson, S. Jha, P. Challis, and R. Kirshner, Harvard-Smithsonian Center for Astrophysics, report that a spectrum of SN 2001df, obtained by P. Berlind on July 19.38 UT with the F. L. Whipple Observatory 1.5-m telescope (+ FAST spectrograph), shows it to be a type-II supernova past maximum light. The spectrum exhibits a flat continuum with P-Cyg lines of hydrogen.

COMETS C/2001 M11, C/2001 N1, C/2001 O1 (SOHO) Further to IAUC 7655, D. Hammer reports measurements of three

additional comets found by M. Oates, X. Leprette, and S. Hönig on SOHO website images. All three objects were visible with the C2 coronagraph, and C/2001 O1 was also visible in C3 images. The reduced astrometry and orbital elements by B. G. Marsden appear on the MPECs cited below; C/2001 N1 does not seem to be a Kreutz sungrazer, whereas the other two comets do.

Comet  2001 UT  α_{2000}  δ_{2000}  MPEC
C/2001 M11 June 27.254  6^{h}18.^{m}6  +22^{o}03′  2001-N24
C/2001 N1  July 5.896  6 55.7  +21 07  2001-N24
C/2001 O1  16.221  7 29.8  +19 15  2001-O03

SUPERNOVA 2001cm IN NGC 5965 CCD R magnitudes obtained by K. Hornoch, Lelekovice, Czech Repub-

lic: June 12.997 UT, 16.2; 15.949, 16.3; 20.998, 16.3; 24.962, 16.4; 25.997, 16.4; 27.925, 16.4; July 12.931, 16.8; 14.901, 16.7. Unfiltered CCD (red) magnitudes by E. Prosperi, Larciano, Italy: June 21.885, 16.7; 24.845, 16.5.

2001 July 19  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2001 CBAT  Daniel W. E. Green